Solve the preblem 6) Suppose the sales of a particular brand of appliance satisfy the relationship So)-50x+5000 where S(x) represents the mumber of sales in year x, with s-0 corresponding to 1982 In wh would the sales be 52502 A) 1990 B) 1985 C) 1988 D) 1987 Find the intercepts and then graph the line. 72x-y-4 A) x-intercept: (2,0x y-intercept (0,4)
Answer #1

Solution : ( 6 )

D) 1987

Explanation :

x = 1987 - 1982 = 5

Now,

S(x) = 50x + 5000

S(5) = 50 * 5 + 5000 = 250 + 5000 = 5250

By this we know that, the sale would be 5250 in the year of 1987

---------------------------------------------------------------------------------------------------------------------------------------------

Solution : ( 7 )

Given equation : 2x -y-4

\mathrm{To\:find\:\:}x\mathrm{-intercept}\::

\mathrm{Put\:\:\:}y=0\:\mathrm{\:\:in\:the\:given\:equation:\:\:\:}2x-y=4

2x - (0) = 4

2x= 4

x=\frac{4}{2}=2

x=2

\mathrm{and}

\mathrm{To\:find\:\:}y\mathrm{-intercept}\::

\mathrm{Put\:\:\:}x=0\:\mathrm{\:\:in\:the\:given\:equation:\:\:\:}2x-y=4

2(0) - y = 4

-y= 4

y=-4

\mathrm{Hence,}

x\mathrm{-intercept=}\left(2,\:0\right),\quad\:y\mathrm{-intercept=}\left(0,\:-4\right)